# Identifying intermediate mass binary black hole mergers in AGN disks using LISA

### v1

@@ Abstract
We show that Laser Interferometer Space Antenna can uniquely identify the sites of intermediate-mass binary black hole (IMBBH) mergers if they occur in Active Galactic Nuclei (AGN) disks with a gas density $\rho\geq 10^{-12}\,{\rm g/cc}$ via measurement of dynamical friction effect in the gravitational waveform. We find that even a single observation of a gravitational wave source with a total mass of $10^{3}M_{\odot}$ and a mass ratio of 2 at a luminosity distance of 3 Gpc is sufficient to confidently associate the merger to be in an AGN disk with a density $\sim 10^{-12}\,{\rm g/cc}$, as it allows estimation of the density with an error bar ${\cal O}(100\%)$. This provides a new way of inferring AGN disk densities that complement traditional X-ray observations. Further, we find that neglecting the presence of environmental effects in the waveform models used for parameter estimation can bias the chirp mass, mass ratio and arrival time of a merger. If not corrected, this can significantly impact our ability to carry out multiband data analysis of IMBBHs that combines information from LISA and the ground-based gravitational wave detectors.
@@ I Introduction
Active Galactic Nuclei disks (AGN disks) are excellent sites for rapid and efficient growth of intermediate mass black holes (IMBHs) [1, 2, 3, 4, 5, 6, 7]. Compared to other astrophysical environments, such as nuclear star clusters (NSC) and globular clusters (GC), the cores of AGN disks can have densities between $10^{-12}-10^{-6}\,{\rm g/cc}$ making them the densest astrophysical environments where such mergers occur [8, 9, 10]. With its capability to detect IMBH binaries at high redshifts with high signal-to-noise ratios (SNRs), the planned Laser Interferometer Space Antenna (LISA) mission [11] could observe a population of intermediate-mass binary black hole (IMBBH) mergers occurring in AGN disks.
@@
It is known that the phase evolution of gravitational wave signals of nonvacuum BBH mergers differs from their vacuum counterparts [14, 15]. The physical processes that can contribute to this dephasing include accretion onto the black hole (BH) [16], dynamical friction due to mass segregation [17, 18, 19, 20], etc. As this dephasing occurs at (effective) negative post-Newtonian (PN) orders (where low-frequency effects dominate in the early part of the inspiral), it is not measurable with the low-frequency sensitivities of the current or proposed future ground-based GW detectors [21, 22]. However, we argue that mergers of IMBBHs with masses in the range $\sim 10^{2}-10^{4}\,M_{\odot}$ with high SNRs, which will be observed by LISA up to redshift of 1 [23, 24, 25], can give us a unique opportunity to precisely measure the density of the medium if the densities are above $\sim 10^{-12}\,{\rm g/cc}$. As this is the typical density of AGN disks [10, 8, 26], and few astrophysical environments can reach such high densities, this implies LISA can confidently identify IMBBH mergers in AGN disks for a subset of the population at low redshifts ($z\leq 0.5$).
@@ I.1 A quick review of past works
References [14, 15] were among the first explorations of the effects of astrophysical environments on gravitational waveforms (see also Ref. [27]). These works studied a suite of physical effects that are related to the interaction of binary with the environment. They concluded that the impact of such environmental effects on inspiral and ringdown signals is generally small enough not to hinder precision inference, except possibly in the case of extreme mass-ratio inspirals [14]. Since then, several studies have investigated the detectability of these environmental effects across different frequency bands of the GW spectrum, covering the detectability of environmental effects in the context of supermassive  [19, 28, 29, 30, 31, 32, 33, 34, 35, 36] intermediate-mass [37, 38, 39, 16] and stellar-mass black holes [40, 41, 42, 43, 22, 21, 44, 16, 45].
@@ II.1 Gravitational waveforms with environmental effect
In presence of astrophysical environments, the dynamics of a compact binary differs from that in vacuum. These environments could be distribution of masses or gas, which causes dynamical friction (DF), accretion of gas onto the BH or the presence of a perturbing third body that can induce anomalous acceleration of the center of mass. Among these effects, DF is expected to induce the most significant observable change in the GW phase for IMBBH mergers in AGN disks with high gas density. Therefore, we focus solely on the effect of DF on the GW phasing of IMBBHs [^1] It is possible that the BHs involved in the merger could be accreting and/or that the binary experiences acceleration of the CM due to the central SMBH. We have checked a few representative cases, and the inclusion of these effects does not change our main results.. Our treatment of the phasing, which accounts for DF, closely follows those of Refs. [62, 17, 18, 63, 21].
@@
In the above equation, $\eta=\frac{q}{(1+q)^{2}}$, where $q$ is the ratio of the primary mass to the secondary mass of the binary, called the symmetric mass ratio, and $M_{c}$ is the source-frame chirp mass, defined as $\rm{M_{c}}=\rm{M}\eta^{3/5}$, with $\rm{M}$ being the binary’s total mass in the source-frame 222 Due to cosmological expansion, the observed detector frame mass is $(1+z)$ times the source-frame mass. Throughout the paper, we quote the source-frame mass of the IMBBHs. Note that while computing the lower cut-off frequency through Eq. (7), the reshifted chirp mass must be used, as that is what is observed by the detector., $z$ is the cosmological redshift of the source. The factor $\gamma_{\rm DF}$ in Eq. 1 is given by $\gamma_{\rm DF}=-247\,\log\left(\tfrac{f}{f_{\rm DF}}\right)-39+304\,\log(20)+38\,\log\left(\tfrac{3125}{8}\right)$, where $f_{\rm DF}=\tfrac{c_{s}}{22\pi M}$, and $c_{s}$ is the speed of sound in the astrophysical environment with density $\rho$. Here, we have assumed $c_{s}=0.01c$, where c is the speed of light.
@@ III.1 Projected error bars on the density of the medium using LISA
In Fig.(1a), we consider IMBBHs with a total mass ranging between $350-6500M_{\odot}$ with a mass ratio $q=2$ and spins aligned with respect to the orbital angular momentum vector, with magnitudes $\chi_{1}=0.2,\chi_{2}=0.1$. We assume the sources to be at a fiducial luminosity distance of 3 Gpc. The chosen mass range ensures that the SNR is always greater than 10 for both LISA and ET, which serves as our representative for the next-generation ground-based observatory. The $1-\sigma$ error bars on the scaled density $(\bar{\rho})$ are converted to the fractional error on the density which is shown in the top plot of Fig.(1a).
For medium with $\rho=10^{-8}\,{\rm g/cc}$ and $10^{-10}\,{\rm g/cc}$, the relative errors are of the order of a few percent, implying that LISA will be able to place very tight constraints on the properties of the medium in these cases. Even for densities of $\sim$ $10^{-12}{\rm g/cc}$, a relative error ${\cal O}(100\%)$ would be very useful in understanding the non-vaccuum environment of the merger and would suffice to identify the merger sites to be AGN disks, which are the most natural candidates to host environments with densities of this order. Multibanding improves these errors by a factor of $2-5$ for those IMBBHs that can be observed in both bands. In the mass range we consider, the best measurement achieves an error bar of $\sim 0.6\%$ for a $1000M_{\odot}$ binary with $q=2$ with multibanding.
Figure (1b) shows the dependence of the errors on the mass ratio for $\rho=10^{-10}$ g/cc. As the mass ratio increases, the fractional errors decrease. This behavior can be attributed to the strong dependence of the dynamical friction term in the phase on the symmetric mass ratio (see Eq. (1)). The non-monotonic nature of the curves in Fig. 1b can be attributed to multibanding of the low- and high- frequency data of LISA and ET, respectively, which gives rise to a minimum around $10^{3}M_{\odot}$.

### v2

@@ Abstract
We show that Laser Interferometer Space Antenna can uniquely identify the sites of intermediate-mass binary black hole (IMBBH) mergers if they occur in Active Galactic Nuclei (AGN) disks with a gas density $\rho\geq 10^{-12}\,{\rm g/cc}$ via measurement of dynamical friction effect in the gravitational waveform. We find that even a single observation of a gravitational wave source with a total mass of $10^{3}M_{\odot}$ and a mass ratio of 2 at a luminosity distance of 3 Gpc is sufficient to confidently associate the merger to be in an AGN disk with a density $\sim 10^{-12}\,{\rm g/cc}$, as it allows estimation of the density with an error bar ${\cal O}(100\%)$. This provides a new way of inferring AGN disk densities that complement traditional X-ray observations. Further, we find that neglecting the presence of environmental effects in the waveform models used for parameter estimation can bias the chirp mass, mass ratio and arrival time of a merger. If not corrected, this can significantly impact our ability to carry out multiband data analysis of IMBBHs that combines information from LISA and ground-based gravitational wave detectors.
@@ I Introduction
Active Galactic Nuclei disks (AGN disks) are excellent sites for rapid and efficient growth of intermediate mass black holes (IMBHs) [1, 2, 3, 4, 5, 6, 7]. Compared to other astrophysical environments, such as nuclear star clusters (NSC), which have densities ranging from $\sim 10^{-18}\rm g/cm^{3}$ to $\sim 10^{-15}\rm g/cm^{3}$ [8], and globular clusters (GC) with densities between $10^{-21}-10^{-18}\,{\rm g/cm^{3}}$ [9], the cores of AGN disks can have densities between $10^{-12}-10^{-6}\,{\rm g/cm^{3}}$ making them the densest astrophysical environments where such mergers occur [10, 11, 12]. With its capability to detect IMBH binaries at high redshifts with high signal-to-noise ratios (SNRs), the planned Laser Interferometer Space Antenna (LISA) mission [13] could observe a population of intermediate-mass binary black hole (IMBBH) mergers occurring in AGN disks.
@@
It is known that the phase evolution of gravitational wave signals of nonvacuum BBH mergers differs from their vacuum counterparts [16, 17]. The physical processes that can contribute to this dephasing include accretion onto the black hole (BH) [18], dynamical friction due to mass segregation [19, 20, 21, 22], etc. As this dephasing occurs at (effective) negative post-Newtonian (PN) orders (where low-frequency effects dominate in the early part of the inspiral), it is not measurable with the low-frequency sensitivities of the current or proposed future ground-based GW detectors [23, 24]. However, we argue that mergers of IMBBHs with masses in the range $\sim 10^{2}-10^{4}\,M_{\odot}$ with high SNRs, which will be observed by LISA up to very high redshifts  [25, 26, 27], can give us a unique opportunity to precisely measure the density of the medium if the densities are above $\sim 10^{-12}\,{\rm g/cc}$. As this is the typical density of AGN disks [12, 10, 28], and few astrophysical environments can reach such high densities, this implies LISA can confidently identify IMBBH mergers in AGN disks for a subset of the population at low redshifts ($z\leq 0.5$).
@@ I.1 A quick review of past works
References [16, 17] were among the first explorations of the effects of astrophysical environments on gravitational waveforms (see also Ref. [29]). These works studied a suite of physical effects that are related to the interaction of binary with the environment. They concluded that the impact of such environmental effects on inspiral and ringdown signals is generally small enough not to hinder precision inference, except possibly in the case of extreme mass-ratio inspirals [16]. Since then, several studies have investigated the detectability of these environmental effects across different frequency bands of the GW spectrum, covering the detectability of environmental effects in the context of supermassive  [21, 30, 31, 32, 33, 34, 35, 36, 37, 38] intermediate-mass [39, 40, 41, 18, 42, 43] and stellar-mass black holes [44, 45, 46, 47, 24, 23, 48, 18, 49].
@@ II.1 Gravitational waveforms with environmental effect
In presence of astrophysical environments, the dynamics of a compact binary differs from that in vacuum. These environments could be distribution of masses or gas, which causes dynamical friction (DF), accretion of gas onto the BH or the presence of a perturbing third body that can induce anomalous acceleration of the center of mass. Among these effects, DF is expected to induce the most significant observable change in the GW phase for IMBBH mergers in AGN disks with high gas density. Therefore, we focus solely on the effect of DF on the GW phasing of IMBBHs [^1] It is possible that the BHs involved in the merger could be accreting and/or that the binary experiences acceleration of the CM due to the central SMBH. We have checked a few representative cases, and the inclusion of these effects does not change our main results. For example, the fractional error on density (see Sec. III.1) of $10^{-10}$ $\rm g/cm^{3}$ remains of the order of $\sim 10^{-3}$ for $\rm M=1500\,M_{\odot},q=5$ binary when computed from only dynamical friction and including both dynamical friction and accretion into the GW phase.. Our treatment of the phasing, which accounts for DF, closely follows those of Refs. [66, 19, 20, 67, 23].
@@
In the above equation, $\eta=\frac{q}{(1+q)^{2}}$, where $q$ is the ratio of the primary mass to the secondary mass of the binary, called the symmetric mass ratio, and $M_{c}$ is the source-frame chirp mass, defined as $\rm{M_{c}}=\rm{M}\eta^{3/5}$, with $\rm{M}$ being the binary’s total mass in the source-frame 222 Due to cosmological expansion, the observed detector frame mass is $(1+z)$ times the source-frame mass. Throughout the paper, we quote the source-frame mass of the IMBBHs. Note that while computing the lower cut-off frequency through Eq. (7), the reshifted chirp mass must be used, as that is what is observed by the detector., $z$ is the cosmological redshift of the source. The factor $\gamma_{\rm DF}$ in Eq. 1 is given by $\gamma_{\rm DF}=-247\,\log\left(\tfrac{f}{f_{\rm DF}}\right)-39+304\,\log(20)+38\,\log\left(\tfrac{3125}{8}\right)$, where $f_{\rm DF}=\tfrac{c_{s}}{22\pi M}$, and $c_{s}$ is the speed of sound in the astrophysical environment with density $\rho$. To get a representative value of speed of sound, following [10], we select $10^{-10}$ g/$\rm cm^{3}$ as characteristic density of AGN disk which corresponds approximately to the speed of sound being 0.01 times the speed of light. Here, we have assumed $c_{s}=0.01c$, where c is the speed of light for all cases.
@@ II.4 Systematic error computation
Given the high SNRs of the events (see below in Sec. IIIA) considered here, we expect our estimates will be representative of the results that will come from a fully Bayesian treatment of the problem. For instance, our Fisher matrix code results are in good agreement with the Bayesian estimates of [67]. However, a dedicated Bayesian study of the problem would help in putting our estimates on a firmer footing. We plan to undertake this as a follow up project.
@@ III.1 Projected error bars on the density of the medium using LISA
In Fig.(1a), we consider IMBBHs with a total mass ranging between $350-6500M_{\odot}$ with a mass ratio $q=2$ and spins aligned with respect to the orbital angular momentum vector, with magnitudes $\chi_{1}=0.2,\chi_{2}=0.1$. We assume the sources to be at a fiducial luminosity distance of 3 Gpc. The chosen mass range ensures that the SNR of binaries in both LISA and ET, which serves as our representative for the next-generation ground-based observatory, is centered around $\sim\mathcal{O}(100)$. Note that the dimensionless density parameter $\bar{\rho}$ was introduced in order to have a well-behaved Fisher matrix. To visualise the results in a more physically meaningful way, we translate the $1-\sigma$ error bars on the scaled density $(\bar{\rho})$ to the fractional error on the density which is shown in the top plot of Fig.(1a).The $\rho$ in the denominator of $\delta\rho/\rho$ corresponds to the particular density being considered.
For medium with $\rho=10^{-8}\,{\rm g/cc}$ and $10^{-10}\,{\rm g/cc}$, the relative errors are of the order of a few percent, implying that LISA will be able to place very tight constraints on the properties of the medium in these cases. Even for densities of $\sim$ $10^{-12}{\rm g/cc}$, a relative error ${\cal O}(100\%)$ would be very useful in understanding the non-vaccuum environment of the merger and would suffice to identify the merger sites to be AGN disks, which are the most natural candidates to host environments with densities of this order. Multibanding improves these errors by a factor of $2-5$ for those IMBBHs that can be observed in both bands. In the mass range we consider, the best measurement achieves an error bar of $\sim 0.6\%$ for a $1000M_{\odot}$ binary with $q=2$ with multibanding. Figure (1b) shows the dependence of the errors on the mass ratio for $\rho=10^{-10}$ g/cc. As the mass ratio increases, the fractional errors decrease. This behavior can be attributed to the strong dependence of the dynamical friction term in the phase on the symmetric mass ratio (see Eq. (1)). The non-monotonic nature of the curves in Fig. 1b can be attributed to multibanding of the low- and high- frequency data of LISA and ET, respectively, which gives rise to a minimum around $10^{3}M_{\odot}$.